# Generative Location Modeling for Spatially Aware Object Insertion

### v1

@@ D.1 Architecture of the Location Model
Our location model is based on a GPT-2 small architecture (Radford et al., 2019), consisting of 12 layers of Transformer blocks (Vaswani et al., 2017), resulting in 117 million parameters. For image and object class encoding, we use a pre-trained ViT-B model (Wightman, 2019; Dosovitskiy, 2021) for the image encoder and a ViT-B CLIP text encoder (Ilharco et al., 2021; Radford et al., 2021) for the object class. The images are converted into 196 embeddings, while the text is transformed into a single embedding, creating a sequence length of 197 when combined. These 197 embeddings are prepended to our location model before predicting the coordinates. To quantize the coordinates, we use 512 bins for both height and width, resulting in a vocabulary size of 514, including the start-of-sequence (SOS) and end-of-sequence (EOS) tokens.
@@ D.2 Training Details
We train the model using the Adam optimizer (Kingma, 2015) for training the location model and Stochastic Gradient Descent (SGD) for DPO training. The model is trained on the PIPE dataset for 30,000 iterations with a learning rate of 1e-4, incorporating a linear warmup over the first 1,000 steps. Subsequently, we fine-tune the model on the OPA dataset for 3,000 steps and perform DPO for an additional 4,000 steps. For batch size 64, the model can be trained on a single Nvidia V100 GPU.
@@ D.3 Sampling from the Locations
Our autoregressive layout model can follow sampling techniques similar to those used for text generation using Large Language Models (LLMs) (Radford et al., 2019). We sample among the top-$k$ probabilities scaled with a temperature of $1.0$. As illustrated in Figure 14, we find that the higher values of $k$ promotes diversity and thus achieves a higher True Positive Rate (TPR), but also increases the False Positive Rate (FPR). For the main experiments, we use $k=50$ during sampling.
Appendix E Human Evaluation
@@
Appendix F Qualitative Results

### v2

@@ D.1 Architecture of the Location Model
Our location model is based on a GPT-2 small architecture [33], consisting of 12 layers of Transformer blocks [41]. For image and object class encoding, we use a pre-trained ViT-B model [44, 6] for the image encoder and a ViT-B CLIP text encoder [15, 34] for the object class. The images are converted into 196 embeddings, while the text is transformed into a single embedding, creating a sequence length of 197 when combined. These 197 embeddings are prepended to our location model before predicting the coordinates. To quantize the coordinates, we use 512 bins for both height and width, resulting in a vocabulary size of 514, including the start-of-sequence (SOS) and end-of-sequence (EOS) tokens.
The location model, including the ViT-B backbone, comprises a total of 411 million parameters, and loading the weights in float16 precision requires just 2.05 GB of VRAM. Running the model with batch size 1 in float16 precision requires an additional 1.93 GB of VRAM for the activations, meaning the model can easily be run on consumer-grade hardware. On a Nvidia Tesla V100 GPU, inference runs in 0.03 seconds. For editing images, we first sample locations, unload the model from the GPU, and perform inpainting, ensuring no additional memory overhead beyond what the inpainting model requires.
In comparison, the PowerPaint inpainting model has three components: a VAE with 83.7 million parameters, a text encoder with 123.1 million parameters, and a UNet with 859.5 million parameters. Although the location model is nearly half the size of the UNet in terms of parameters, it is used only once in advance (similar to the VAE and text encoder), whereas the UNet is used at every reverse step. On the same GPU, the it takes 7.10 seconds to perform 50 reverse steps, meaning the inference overhead of the location model is 0.4%, an acceptable cost for the improvement in generation quality.
@@ D.2 Training Details
We train the model using the Adam optimizer [19] for training the location model and Stochastic Gradient Descent (SGD) for DPO training. The model is trained on the PIPE dataset for 30,000 iterations with a learning rate of 1e-4, incorporating a linear warmup over the first 1,000 steps. Subsequently, we fine-tune the model on the OPA dataset for 3,600 steps and perform DPO for an additional 4,600 steps. For batch size 64, the model can be trained on a single Nvidia V100 GPU.
@@ D.3 Sampling from the Locations
Our autoregressive layout model can follow sampling techniques similar to those used for text generation using Large Language Models (LLMs) [33]. We sample among the top-$k$ probabilities scaled with a temperature of $1.0$. As illustrated in Fig. 12, we find that the higher values of $k$ promotes diversity and thus achieves a higher True Positive Rate (TPR), but also increases the False Positive Rate (FPR). For the main experiments, we use $k=50$ during sampling.
Appendix E Diversity of the Sampled Location
An essential aspect of location models is their ability to identify diverse plausible locations, as multiple potential placements often exist for a given object. To quantify this diversity, we apply Non-maximum Suppression (NMS) to the predicted locations and count the remaining bounding boxes. Specifically, we perform a standard NMS with a threshold of 0.7 on a set of 100 predicted locations for the same object in the same scene. We observe that 82.5 bounding boxes remain after performing NMS, which highlights the diversity of our location model.
Appendix F Human Evaluation
@@
Appendix G Qualitative Results
@@
Additionally, we show qualitative samples highlighting generalization capability in Fig. 16. As the class is encoded using a CLIP text encoder, we can condition on class labels that are not present in the data and expect a reasonable output if the corresponding CLIP embedding is sufficiently close to existing categories. We show that we can add another instance of an already present object without our bounding box prediction coinciding with the existing instance, and that we can add instances from other classes in OPA and out-of-domain classes.
Figure 16: Qualitative examples showcasing generalization ability. We ask our model to add one more instance of an already present class (repeated), an instance of a different class (in-domain), and an instance of a class that does not exist in the OPA dataset (unseen).